OpenStudy (anonymous):

A firecracker shoots up from a hill 145 feet high, with an initial speed of 80 feet per second. Using the formula H(t) = -16t2 + vt + s, approximately how long will it take the firecracker to hit the ground? 13 seconds 11 seconds Nine seconds Six seconds

OpenStudy (anonymous):

@sleepyjess @amistre64

OpenStudy (astrophysics):

\[h(t) = -16t^2+vt+s\] we are given the initial speed to be 80 feet/s with a height of 145 feet high, so lets plug that in. \[h(t) = -16t^2+80t+145\] for it to hit the ground we have to solve for h(t) = 0, so we can use the quadratic formula, I leave that to you.\[-16t^2+80t+145=0\] where the quadratic formula is\[ t = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] as we're solving for time.

OpenStudy (astrophysics):

Also note you want a positive value.

OpenStudy (anonymous):

Thank You

OpenStudy (astrophysics):

Np :)

OpenStudy (anonymous):

Is it D?

OpenStudy (astrophysics):

Yes, that sounds pretty good to me :)